# Hilbert schemes on blowing ups and the free Boson

### v2

@@ 3.3. Semi-orthogonal decomposition and decomposition of the diagonal
Moreover, Jiang [13] proved that the functors are given as Fourier-Mukai transforms through the incidence varieties. Now we introduce a special case when the rank of $\mathcal{E}$ is $1$:
@@
Let $\Psi_{d,d-1}^{R}$ be the right adjoint functor of $\Phi_{d,d-1}$ and $\Psi_{d,d}^{L}$ be the left-adjoint fucntor of $\Phi_{d,d}$. Then there exsits a canonical triangle of Fourier-Mukai transforms
(3.3) $\Phi_{d,d-1}\Psi_{d,d-1}^{R}\to\mathcal{O}_{\Delta}\to\Phi_{d,d}\Psi_{d,d}^{L}$
@@
(3.4) $\Psi_{d,d-1}^{R}\Phi_{d,d-1}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d}^{L}\Phi_{d,d}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d-1}^{R}\Phi_{d,d}\cong 0,\quad\Psi_{d,d}^{L}\Phi_{d,d-1}\cong 0.$
@@
We only need to show natural transforms of functors are induced by morphisms of Fourier-Mukai transforms. As all the setting can be naturally lifted to $\infty$-categories, it follows from Theorem 1.2(2) of [3]. ∎
4. Mukai pairing and cohomological correspondences
In this section we always assume that $X$ and $Y$ be smooth projective varieties. Given $f\in Perf(X\times Y)$, it naturally induces a Fourier-Mukai transform
$\Gamma_{X\to Y}^{\mu}:Perf(X)\to Perf(Y),\quad\Gamma_{X\to Y}^{\mu}(\cdot)=\pi_{Y,*}(\pi_{X}^{*}(\cdot)\otimes\mu).$
@@ 4. Mukai pairing and cohomological correspondences
One may expect that two transforms can be naturally compared by the Chern chracter, i.e. we have a commutative diagram
(4.1) ${Perf(X)}$${Perf(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{ch}$$\scriptstyle{ch}$$\scriptstyle{\gamma_{X\to Y}^{ch(\mu)}}$
where $ch$ is the Chern chracter. Unfortunately, 4.1 in general is not true due to the Grothendieck-Riemann-Roch theorem. Caldararu [4] found the correct way to modify the above diagram by generalizing the Mukai vectors, which we will recall in this section.
4.1. Fourier-Mukai transforms and cohomological correspondences
@@
The directed Mukai vector of an element $\mu\in Perf(X\times Y)$ is defined by
@@
By taking the first space to be a point we obtain the definition of the Mukai vector of an object $\alpha\in Perf(X)$ as
@@
We define a map $(-)_{*}$ from the Fourier-Mukai transforms on $Perf(X\times Y)$ to cohomological correspondences $\mathbb{H}^{*}(X\times Y,\mathbb{Q})$ by
@@
The map $(-)_{*}$ has the following properties:
@@
    Given Fourier-Mukai transforms $\mu:Perf(X)\to Perf(Y)$, $\nu:Perf(Y)\to Perf(Z)$, we have
@@
    Given a triangle of Fourier-Mukai transforms
    $f\to g\to h\to f[1],$
    we have $g_{*}=f_{*}+h_{*}$.
  4. (4)
@@
    (4.2) ${Perf(X)}$${Perf(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{v}$$\scriptstyle{v}$$\scriptstyle{\gamma_{X\to Y}^{v(\mu)}}$
@@ 4.2. Cohomological correspondences of derived Grassmannians
are (classical) smooth varieties. We apply the Caldararu’s construction for $\Phi_{d_{+},d_{-}}$ and by 3.3 and 3.4 and have
@@
The cohomological correspondences $\Phi_{d,d-1*},\Psi_{d,s-1*}^{R},\Phi_{d,d*},\Psi_{d,d*}^{L}$ satisfy the relations 2.2, if we denote
$E:=\Phi_{d,d-1*},\quad F:=\Psi_{d,d-1*}^{R},\quad K_{1}:=\Phi_{d,d*},\quad K_{-1}:=\Phi_{d,d*}^{L}.$
Now we consider the (co)homological degree of operators
$\Phi_{d,d-1*},\Psi_{d,s-1*}^{R},\Phi_{d,d*},\Psi_{d,d*}^{L}.$
For the Hochschild homology, their homological degree are all 00. But for the cohomology groups and Chow groups, their cohomological degree is not a constant and we need to make a modification. Recall that the incidence variety
@@
The operators $\phi_{d,d-1}$ and $\psi_{d,d-1}$ presevered the cohomological degree and the operators $\phi_{d,d}$ and $\psi_{d,d}$ shifted the Chow group (resp. cohomological, Hodge) degree by $d$ (resp. $2d$, $(d,d)$) and $-d$ (resp. $-2d$, $(-d,-d)$). Moreover, the cohomological correspondences $\phi_{d,d-1},\psi_{d,s-1},\phi_{d,d},\psi_{d,d}$ also satisfy the relations 2.2 if we denote
@@
First we notice that all the constant terms of roots of Todd class and Chern classes of line bundles are $1$. By the virtual Grothendieck-Riemann-Roch theorem (i.e. Theorem 3.23 and Corollary 3.25 of [14]), we also have
@@
where $C_{d_{+},d_{-}}$ have higher degrees than $Vir_{d_{+},d_{-}}$. Thus we have $\Phi_{d,d*}-\phi_{d,d}$ has higher cohomological degree than $\phi_{d,d}$. Similar computations also holds for $\Phi_{d,d-1*}-\phi_{d,d-1},\Psi_{d,d*}^{L}-\psi_{d,d}$ and $\Psi_{d,d*}-\psi_{d,d-1}$. Thus by checking the cohomological grading, the relations 2.2 for $\phi_{d,d-1},\psi_{d,s-1},\phi_{d,d},\psi_{d,d}$ follows from it for $\Phi_{d,d-1*},\Psi_{d,d-1*}^{R},\Phi_{d,d*},\Psi_{d,d*}^{L}$. ∎
@@ 5. Perverse Coherent Sheaves and their moduli spaces
Let $S$ be a smooth surface with a fixed ample line bundle $H$. We denote by
@@ 5.1. Moduli space of perverse coherent sheaves
where $[C]\in H^{1,1}(\hat{S})\cap H^{2}(\hat{S},\mathbb{Q})$ and $[C]^{2}=-1$.
@@ 5.2. Moduli spaces as derived Grassmannians
We first recall cohomological properties of stable coherent sheaves
@@ 6. The admissible $\mathcal{E}$-representation on perverse coherent sheaves
Given a smooth projective surface $S$ and a closed point $o\in S$ and a cohomology theory $\mathbb{H}$, we consider
@@
Other than the grading by the first Chern characters (or equivalently $m$-stability), we also have two gradings: the $q$-degree by associating $\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})$ with degree $n$, the $t$-degree (or bi-degree) which is the cohomology degree. In this section, we prove that
@@
There is a $\mathcal{E}$-representation on $\mathbb{H}_{0}^{S,o}$ such that
@@
    the representation is graded admissible;
@@
    the cohomology degree of $E,F$ are preserved, and the cohomology degree (with respect to Hochschild homology, Hodge cohomolgy and Chow group) of $K_{\pm 1}$ are $0,\pm(l,l),\pm l$ on $\mathbb{H}^{S,o}_{l}$.
@@
and the action of the infinite dimensional Clifford algebra on $F$ is given by geometric correspondences. The cohomological degree (with respect to Hoschild homology, Hodge cohomolgy and Chow group) of a monomial in $F$:
@@
is $0,(x_{i_{1}}+\cdots+x_{i_{n}}-\frac{n^{2}+n}{2},x_{i_{1}}+\cdots+x_{i_{n}}-\frac{n^{2}+n}{2}),x_{i_{1}}+\cdots+x_{i_{n}}-\frac{n^{2}+n}{2}$ respectively and the $q$-degree is equal to the $t$-degree of the Chow group.
@@
In particular, the Heisenberg algebra action on $B$ is induced by geometric correspondences. For a monomial in the Bosonic Fock space
$y_{1}^{i_{1}}y_{2}^{i_{2}}\cdots y_{n}^{i_{n}}$
its cohomological degree with respect to Hoschild homology, Hodge cohomology and Chow group is
$0,\quad\sum_{j=1}^{n}ji_{j}(1,1),\quad\sum_{j=1}^{n}ji_{j}.$
and the $q$-degree is same as the $t$-degree of the Chow group.
@@ 6.1. The geometric correspondences
Then they all acts on $\mathbb{H}_{l,n}^{S,o}$ (for the Hochschild homology, we replace it by $\Psi_{*,*}$ and $\Phi_{*,*}$ respectively).
@@ Appendix A Some algebra computations
In this section, we give the proof of Proposition 2.2, Proposition 2.16, Proposition 2.17 and Proposition 2.19, which involve explicit computations.
@@
and similar equations for $Q(z)$. Moreover, we have
@@
A.2. The proof of Proposition 2.16
Now let $W$ be an graded adimissible $\mathcal{C}$-module. To prove Proposition 2.16, we first prove the following lemmas
Lemma A.2.
Let $\beta$ be a degree 00 operator on $W$, such that $\beta|_{W_{0}}=0$ and $[Q(z),\beta]=0$, i.e.
$[Q_{l},\beta]=0$
for any $l>0$, then $\beta=0$.
Proof.
We prove that $\beta|_{W_{i}}=0$ by the induction on $i$. Assuming that $\beta|_{W_{i}}=0$, then on $W_{i+1}$,
$Q_{l}\beta|_{W_{i+1}}=0$
for all $l>0$ and hence $\beta=0$ by the admissible property of $W$. ∎
Corollary A.3.
Let $\alpha$ and $\beta$ be two degree 00 operators on $W$ such that $\beta|_{W_{0}}=\alpha|_{W_{0}}$ and
$[Q_{l},\beta]=[Q_{l},\alpha]$
for all $l>0$. Then $\alpha=\beta$.
Recalling
@@
We have $\Lambda|_{W_{i}}=i\cdot id$.
@@
for all $l>0$. Hence they are the same by Corollary A.3.
@@
We have
$Q(z)K_{1}=zK_{1}Q(z),\quad K_{-1}P(z)=zP(z)K_{-1},$
i.e. if $j>0$
$Q_{j}K_{1}=K_{1}Q_{j-1},\quad K_{-1}P_{j}=P_{j-1}K_{-1},$
@@
$(\Lambda+id)Q_{j}K_{1}=Q_{j}\Lambda K_{1}=\sum_{l=1}^{\infty}Q_{j}P_{l+1}K_{1}Q_{l}=-\sum_{j=1}^{\infty}P_{l+1}(Q_{j}K_{1})Q_{l}+K_{1}Q_{j-1}$
$\Lambda K_{1}Q_{j-1}=\sum_{l=1}^{\infty}P_{l+1}K_{1}Q_{l}Q_{j-1}=-\sum_{j=1}^{\infty}P_{l+1}(K_{1}Q_{j-1})Q_{j-1}$
Thus $\alpha:=Q_{j}K_{1}-K_{1}Q_{j}$ satisfy the equation that
@@
By induction, the solutions of A.5 is decided by $\alpha|_{W_{0}}$, and hence $\alpha=0$ follows from $\alpha|_{W_{0}}=0$. ∎
@@
The second formula follows from A.6 and Lemma A.2. For the first formula, let $\beta:=K_{1}K_{-1}-Q_{1}P_{1}$. We have
$\beta|_{W_{0}}=0,\quad Q_{1}\beta=0,\quad[Q_{l},\beta]=0$
for $l>1$. Then by induction and the similar method with Lemma A.2, we also have $\beta=0$. ∎
Proof of Proposition 2.16.
@@
Let $H_{\pm k}$, $k\in\mathbb{Z}-\{0\}$ be the generators of the infinite dimensional algebra. By induction we have
Lemma A.8.
For any integer $i\geq 0$, we have
(A.7) $H_{-k}H_{k}^{i}=H_{k}^{i}H_{-k}-ik,\quad H_{-k}^{i}H_{k}=H_{k}H_{-k}^{i}-ik.$
@@
By A.7 we have
@@
By A.8, we have
$(Id-H_{k}\mathcal{L}_{k})W_{k}\subset W_{k-1}.$
Now we consider
$W:=\bigoplus_{i\geq 0}W_{i}.$
and define degree 00 operators
$K_{1}=\bigoplus_{i\geq 0}H_{k},\quad K_{-1}=\bigoplus_{i\geq 0}\mathcal{L}_{k}$
where $H_{0}=\mathcal{L}_{0}:=id$.
The inclusion of $W_{k}\subset W_{k+1}$ naturally induces a degree $1$ operator on $W$:
$E:=\bigoplus_{k\geq 0}id.$
By A.8, the following degree $-1$ operator is well defined
$F:=\bigoplus_{k\geq 0}(Id-H_{k}\mathcal{L}_{k})$
@@
We need to prove that under the above construction, the vector space $W$ forms an admissible graded $\mathcal{E}$-algebra representation.
First we notice that for any $k>0$, $\mathcal{L}_{k}|_{W_{k-1}}=0$ and thus $K_{-1}E=0$ and $FE=id$. The formula $K_{1}K_{-1}+EF=0$ follows from the definition of $K_{\pm 1}$ and $E,F$.
On the other hand, by A.8 we have
@@
Thus $W$ is a graded $\mathcal{E}$-representation and the admissibility of $W$ follows from the admissibility of $\mathcal{W}_{\infty}$. ∎
@@
Let $W$ be an admissible graded $\mathcal{E}$-representation. As the limit of 2.3, on $W_{l}$ ($l>0$) we have
@@
We assume that $|j|\leq|k|$. We prove it by induction on $i$. It is trivial when $i=0$. Assuming it is true for $i-1$, then if $k=\pm i$, A.10 follows from Lemma A.10 and Lemma A.11. Otherwise, by Lemma 2.1 we have

### v3

@@ 3.3. Semi-orthogonal decomposition and decomposition of the diagonal
Moreover, Jiang [14] proved that the functors are given as Fourier-Mukai transforms through the incidence varieties. Now we introduce a special case when $e=1$.
@@
Let $\Psi_{d,d-1}$ be the right adjoint functor of $\Phi_{d,d-1}$ and $\Psi_{d,d}^{L}$ be the left-adjoint fucntor of $\Phi_{d,d}$. Then there exists a canonical triangle of Fourier-Mukai transforms
$\Phi_{d,d-1}\Psi_{d,d-1}\to\mathcal{O}_{\Delta}\to\Phi_{d,d}\Psi_{d,d}^{L}$
@@
$\Psi_{d,d-1}\Phi_{d,d-1}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d}^{L}\Phi_{d,d}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d-1}\Phi_{d,d}\cong 0,\quad\Psi_{d,d}^{L}\Phi_{d,d-1}\cong 0.$
@@
trans We only need to show natural transformations of functors are induced by morphisms of Fourier-Mukai kernels. As all the settings can be naturally lifted to $\infty$-categories, it follows from Theorem 1.2(2) of [3]. ∎
From now on, we denote
$\Psi_{d,d-1}:=\Psi_{d,d-1}^{R},\quad\Psi_{d,d}:=\Psi_{d,d}^{L}.$
Corollary 3.9.
Let $[\Psi_{d,d-1}]$, $[\Psi_{d,d}]$, $[\Psi_{d,d-1}]$, $[\Psi_{d,d}]$ be the corresponding correspondences at the level of Grothendieck group of perfect complexes. Then we have
(3.2) $[\Psi_{d,d-1}][\Phi_{d,d-1}]=id=[\Psi_{d,d}][\Phi_{d,d}]\quad[\Psi_{d,d-1}][\Phi_{d,d}]=0=[\Psi_{d,d}][\Phi_{d,d-1}],$
and
(3.3) $[\Phi_{d,d-1}][\Psi_{d,d-1}]+[\Phi_{d,d}][\Psi_{d,d}]=id.$
4. Mukai pairing, K-theoretic and cohomological correspondences
Let $\mathbb{H}$ be a cohomology theory other than the Grothendieck groups of coherent sheaves. In this section, we always assume that $X$ and $Y$ are smooth projective varieties. Given an element in the Grothendieck group of coherent sheaves $f\in K(X\times Y)$, one can define the K-theoretic correspondence
$\Gamma_{X\to Y}^{\mu}:K(X)\to K(Y),\quad\Gamma_{X\to Y}^{\mu}(\cdot)=\pi_{Y,*}(\pi_{X}^{*}(\cdot)\otimes\mu).$
@@ 4. Mukai pairing, K-theoretic and cohomological correspondences
One may expect that two transforms can be naturally compared by the Chern character, i.e. we have a commutative diagram
(4.1) ${K(X)}$${K(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{ch}$$\scriptstyle{ch}$$\scriptstyle{\gamma_{X\to Y}^{ch(\mu)}}$
where $ch$ is the Chern character. Unfortunately, 4.1 in general is not true due to the Grothendieck-Riemann-Roch theorem. Caldararu [4] found the correct way to modify the above diagram by generalizing the Mukai vectors, which we will recall in this section.
4.1. K-theoretic and cohomological correspondences
@@
The directed Mukai vector of an element $\mu\in K(X\times Y)$ is defined by
@@
By taking the first space to be a point we obtain the definition of the Mukai vector of an object $\alpha\in K(X)$ as
@@
We define a map $(-)_{*}$ from the $K$-theoretic correspondences on $K(X\times Y)$ to cohomological correspondences $\mathbb{H}^{*}(X\times Y,\mathbb{Q})$ by
@@
The map $(-)_{*}$ is linear with the following properties:
@@
    Given correspondences $\mu:K(X)\to K(Y)$, $\nu:K(Y)\to K(Z)$, we have
@@
    (4.2) ${K(X)}$${K(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{v}$$\scriptstyle{v}$$\scriptstyle{\gamma_{X\to Y}^{v(\mu)}}$
@@ 4.2. Cohomological correspondences of derived Grassmannians
are (classical) smooth varieties. We apply the Caldararu’s construction for $[\Phi_{d_{+},d_{-}}]$ and by 3.2 and 3.3 and have
@@
The cohomological correspondences $[\Phi_{d,d-1}]_{*},[\Psi_{d,{d-1}}]_{*},[\Phi_{d,d}]_{*},[\Psi_{d,d}]_{*}$ satisfy the relations 2.1, if we denote
$E:=[\Phi_{d,d-1}]_{*},\quad F:=[\Psi_{d,d-1}]_{*},\quad K_{1}:=[\Phi_{d,d}]_{*},\quad K_{-1}:=[\Phi_{d,d}]_{*}.$
The correspondences in Corollary 4.4 are homogeneous with degree 00 for the Hochschild homology, but not for the singular cohomology or Chow groups and we need to modify them. Recall that the incidence variety
@@
The operators $\phi_{d,d-1}$ and $\psi_{d,d-1}$ preserved the cohomological degree and the operators $\phi_{d,d}$ and $\psi_{d,d}$ shifted the Chow group (resp. cohomological, Hodge) degree by $d$ (resp. $2d$, $(d,d)$) and $-d$ (resp. $-2d$, $(-d,-d)$). Moreover, the cohomological correspondences $\phi_{d,d-1},\psi_{d,s-1},\phi_{d,d},\psi_{d,d}$ also satisfy the relations 2.1 if we denote
@@
First, we notice that all the constant terms of roots of the Todd class and Chern classes of line bundles are $1$. By the virtual Grothendieck-Riemann-Roch theorem (i.e. Theorem 3.23 and Corollary 3.25 of [15]), we also have
@@
where $C_{d_{+},d_{-}}$ have higher degrees than $Vir_{d_{+},d_{-}}$. Thus we have $[\Phi_{d,d}]_{*}-\phi_{d,d}$ has a higher cohomological degree than $\phi_{d,d}$. Similar computations also holds for $[\Phi_{d,d-1}]_{*}-\phi_{d,d-1},[\Psi_{d,d}]_{*}-\psi_{d,d}$ and $[\Psi_{d,d-1}]_{*}-\psi_{d,d-1}$. Thus by checking the cohomological grading, the relations 2.1 for $\phi_{d,d-1},\psi_{d,d-1},\phi_{d,d},\psi_{d,d}$ follows from Corollary 4.4. ∎
@@ 5. Perverse Coherent Sheaves and their moduli spaces
Let $S$ be a smooth projective surface. We denote by
@@ 5.1. Moduli space of perverse coherent sheaves
where $[C]\in H^{1,1}(\hat{S})\cap H^{2}(\hat{S},\mathbb{Q})$ is the fundamental class of $C$ satisfying $[C]^{2}=-1$.
@@ 5.2. Moduli spaces as derived Grassmannians
We first recall the cohomological properties of stable coherent sheaves
@@ 6. The admissible $\mathcal{E}$-representation on perverse coherent sheaves
Given a smooth projective surface $S$ and a closed point $o\in S$ and a cohomology theory $\mathbb{H}$, we consider the grading which we call as the $l$-grading
@@
Other than the $l$-grading, we also have two grading: the $q$-degree by associating $\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})$ with degree $n$, the $t$-degree (or bi-degree) which is the cohomology degree. We denote
$1_{\mathbb{H}}:=\begin{cases}0,\text{ if }\mathbb{H}\text{ is the Hochschild homology group},\\ 1,\text{ if }\mathbb{H}\text{ is the Chow group or Borel-Moore homology group},\\ (1,1)\text{ if }\mathbb{H}\text{ is the Hodge cohomology group.}\\ \text{Undefined}\text{ if }\mathbb{H}\text{ is the Grothendieck group of coherent sheaves.}\end{cases}$
In this section, we prove that
@@
There is a $\mathcal{E}$-representation on $\mathbb{H}^{S,o}$ such that
@@
    Under the $l$-grading, the representation is graded admissible;
@@
    the $t$-degree of $E,F$ are preserved, and the $t$-degree $K_{\pm 1}$ is ±$\pm l\cdot 1_{\mathbb{H}}$ on $\mathbb{H}^{S,o}_{l}$.
@@
and the action of the infinite-dimensional Clifford algebra on $F$ is given by geometric correspondences. The $t$-degree of a monomial in $F$:
@@
is $(i_{1}+\cdots+i_{n}-\frac{n^{2}+n}{2})\cdot 1_{\mathbb{H}}$ and the $q$-degree is $i_{1}+\cdots+i_{n}-\frac{n^{2}+n}{2}$.
@@
In particular, the Heisenberg algebra action on $B$ is induced by geometric correspondences. For a monomial $y_{\lambda}$ in the Bosonic Fock space, its $t$-degree is $|\lambda|\cdot 1_{\mathbb{H}}$ and the $q$-degree is $|\lambda|$.
@@ 6.1. The geometric correspondences
Then they all act on $\mathbb{H}_{l,n}^{S,o}$ (for the Grothendieck Grothendieck or Hochschild homology, we replace it by $\Psi_{*,*}$ and $\Phi_{*,*}$ or $[\Psi_{*,*}]$ and $[\Phi_{*,*}]$ respectively).
@@ Appendix A Some algebra computations
In this section, we give the proof of Proposition 2.2, Proposition A.4, Proposition 2.17 and Proposition 2.19, which involve explicit computations.
@@
Moreover, we have
@@
A.2. The $\mathcal{E}$-representation on a graded admissible $\mathcal{C}$-module
Let $W$ be a graded admissible $\mathcal{C}$-module. We formally define
For any operator $T$ on $W$, we consider
$\Gamma(T)(z):=\sum_{i\in\mathbb{Z}}\Gamma_{i}(T)z^{i}$
such that
which is well-defined due to the admissible property. We also formally written
$\Gamma(T)(z):=P(z)TQ(z^{-1}).$
We denote
@@ A.2. The $\mathcal{E}$-representation on a graded admissible $\mathcal{C}$-module
We have $\Lambda|_{W_{i}}=i\cdot id$, which is an invertible operator when $i>0$.
To prove Lemma A.2, we first prove that
Lemma A.3.
Let $\alpha$ and $\beta$ be two degree 00 operators on $W$. Then $\alpha=\beta$ if
$\beta|_{W_{0}}=\alpha|_{W_{0}}\quad\text{ and }\quad[Q_{l},\beta]=[Q_{l},\alpha]\text{ for all }l>0.$
@@
We prove that $\alpha-\beta=0$ on all $W_{i}$ by the induction on $i$. Assuming that $\alpha-\beta|_{W_{i}}=0$, then on $W_{i+1}$,
$Q_{l}(\alpha-\beta)|_{W_{i+1}}=(\alpha-\beta)Q_{l}|_{W_{i}}=0\text{ for all }l>0.$
Hence $\alpha-\beta=0$ on $W_{i+1}$ by the admissible property of $W$. ∎
Proof of Lemma A.2.
@@
for all $l>0$. Hence they are the same by Lemma A.3. ∎
  1. (1)
    $K_{1},K_{-1}$ are identity on $W_{0}$;
  2. (2)
    we have
    $\Lambda K_{1}=\Gamma_{1}(K_{1}),\quad\Lambda K_{-1}=\Gamma_{-1}(K_{-1}).$
We also define
$E:=P_{1}K_{1},\quad F:=K_{-1}Q_{1}.$
Proposition A.4.
The operators $K_{1},K_{-1},E,F$ induce a graded admissible $\mathcal{E}$-module on $W$.
To prove Proposition A.4, we first prove the following lemmas
@@
The operators $K_{1},K_{-1},E,F$ satisfy the equation A.4, i.e.
$Q_{j}K_{1}=K_{1}Q_{j-1},\quad K_{-1}P_{j}=P_{j-1}K_{-1},\text{ if }j>0$
$P_{j}K_{1}=K_{1}P_{j-1},\quad K_{-1}Q_{j}=Q_{j-1}K_{-1},\text{ if }j>1,$
@@
$(\Lambda+id)Q_{j}K_{1}=Q_{j}\Lambda K_{1}=\sum_{l=1}^{\infty}Q_{j}P_{l+1}K_{1}Q_{l}=-\sum_{l=1}^{\infty}P_{l+1}(Q_{j}K_{1})Q_{l}+K_{1}Q_{j-1}$
$\Lambda K_{1}Q_{j-1}=\sum_{l=1}^{\infty}P_{l+1}K_{1}Q_{l}Q_{j-1}=-\sum_{l=1}^{\infty}P_{l+1}(K_{1}Q_{j-1})Q_{l}$
Thus $\alpha:=Q_{j}K_{1}-K_{1}Q_{j-1}$ satisfy the equation that
@@
By induction, the solutions of A.7 is decided by $\alpha|_{W_{0}}$, and hence $\alpha=0$ follows from $\alpha|_{W_{0}}=0$. ∎
@@
The second formula follows from A.8 and Lemma A.3. For the first formula, let $\beta:=K_{1}K_{-1}-Q_{1}P_{1}$. We have
$\beta|_{W_{0}}=0,\quad Q_{1}\beta=0,\text{ and }\quad[Q_{l},\beta]=0\text{ for all }l>1.$
Then by induction and the similar method with Lemma A.3, we also have $\beta=0$. ∎
Proof of Proposition A.4.
@@
Let $H_{\pm k}$, $k\in\mathbb{Z}-\{0\}$ be the generators of the infinite-dimensional Heisenberg algebra.
@@
By induction, for any integer $i\geq 0$, we have
(A.10) $H_{-k}H_{k}^{i}=H_{k}^{i}H_{-k}-ik,\quad H_{-k}^{i}H_{k}=H_{k}H_{-k}^{i}-ik.$
By A.10 we have
@@
We only need to prove that it forms a graded $\mathcal{E}$-representation, and the admissibility follows from the admissibility of $\mathcal{W}_{\infty}$.
First, we notice that for any $k>0$, the restriction of
$\mathcal{L}_{k}:W_{k}\to W_{k}$
at $W_{k-1}$ is zero. Thus we have
$K_{-1}E=0,\quad FE=id.$
The formula $K_{1}K_{-1}+EF=0$ follows from the definition of $K_{\pm 1}$ and $E,F$.
On the other hand, by A.9 we have
@@
Thus $W$ is a graded $\mathcal{E}$-representation. ∎
@@
Let $W$ be an admissible graded $\mathcal{E}$-representation. As the limit of 2.2, on $W_{l}$ ($l>0$) we have
@@
We assume that $|j|\leq|k|$. We prove it by induction on $i$. It is trivial when $i=0$. Assuming it is true for $i-1$, then if $k=\pm i$, A.12 follows from Lemma A.9 and Lemma A.10. Otherwise, by Lemma 2.1 we have